OpenStudy (anonymous):

Find the 7th term of the geometric sequence with the given terms. a4 = -8 and a5 = -40; a7 = _______.

OpenStudy (lgbasallote):

common ratio = -40/-8 = 5 now use \[a_n = a_1 r^{n-1}\] \[a_7 = a_4 r^{4-1}\] \[a_7 = -8 \times 5^{3}\] solve for that